OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

I was able to determine by using the formula r^2=x^2+y^2 r=2 times the square root of 2

OpenStudy (anonymous):

Also by using the formula tan theta=y/x Then theta=(-pi/4)+npi

OpenStudy (anonymous):

For my first comment its actually r=plus/minus 2 times the square root of 2

ganeshie8 (ganeshie8):

looks good^

ganeshie8 (ganeshie8):

All the representations : \(\large \left( 2\sqrt{2}, ~~\frac{-\pi}{4} + 2n\pi\right)\) and \(\large \left( -2\sqrt{2}, ~~\frac{-\pi}{4}+\pi + 2n\pi\right)\)

OpenStudy (anonymous):

Which one of these would be the answer then? (2 square root of 2, 135°), (-2 square root of 2, 315°) (2 square root of 2, 315°), (-2 square root of 2, 135°)

ganeshie8 (ganeshie8):

plugin n=1 in first solution plugin n=0 in second solution

OpenStudy (anonymous):

so it would be the second on then

ganeshie8 (ganeshie8):

\(\large \left( 2\sqrt{2}, ~~\frac{-\pi}{4} + 2\pi\right)\) and \(\large \left( -2\sqrt{2}, ~~\frac{-\pi}{4}+\pi +0\right)\)

OpenStudy (anonymous):

(2 square root of 2, 315°), (-2 square root of 2, 135°)

ganeshie8 (ganeshie8):

Yep !